OpenStudy (anonymous):

Find an appropriate viewing window for the function, using the given x-values. y=x^2 +20, for x-values between -6 and 6 Choose the most appropriate viewing window boundaries for the function. a) [-6,6] by [-10,10] b) [-6,6] by [20,56] c) [-6,6] by [0,56] d) [-6,6] by [0,20]

OpenStudy (anonymous):

you need to think about what the minimum value for the function would be. Where would the function be at its minimum? Another thing you will need to check is the value of the function at x = 6 (as -6 would give the same value with it having an x squared)

OpenStudy (anonymous):

hmmm d?

OpenStudy (anonymous):

if you put x = 6, what do you get for y? Also, if it is a parabola it will have a minimum at the turning point (due to it being positive)

OpenStudy (anonymous):

56

OpenStudy (anonymous):

yup, so you know that the graph has to go up to 56. now if you find the smallest value you will have the correct range. Hint: what is the smallest that x squared can be?

OpenStudy (anonymous):

and use this in the formula to find the smallest that y can be

OpenStudy (anonymous):

So it's c

OpenStudy (anonymous):

well that is the smallest that x can be but if x =0, what would y be?

OpenStudy (anonymous):

that was probably my bad for not putting the previous 2 posts together

OpenStudy (anonymous):

Pardon?

OpenStudy (anonymous):

we know that the smallest that x can be is 0, so xsquared is 0. putting this into the equation gives: y = 0 + 20 = 20 That would mean the smallest the equation could be is 20 and the highest in the range was 56 i.e. b

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

that make sense?

OpenStudy (anonymous):

yeah, what does i.e. mean?

OpenStudy (anonymous):

in example. basically i just meant it was B

OpenStudy (anonymous):

How would I graph this?

OpenStudy (anonymous):

well you know it is a parabola and that it is positive ("u" shaped) then you just put in the points that you know: x=0, y=20 and x=-6, y=56 and x=6, y=56 see if you can draw it and i will post a pic if needs be

OpenStudy (anonymous):

like?

OpenStudy (anonymous):

yup just make sure you include the coordinates of the points and thats you done

OpenStudy (anonymous):

You've been a big help with this problem. Thank-you

OpenStudy (anonymous):

no problem